The sum of the measures of angle X and angle Y is 90. If the measure of angle X is 30 less than twice the measure of angle Y, what is the measure of angle X?
Answer by ReadingBoosters(3246)

Establish equations
x+y = 90
x = 2y - 30
...
Substitute the x = into the first equation
2y - 30 + y = 90
...
solve for y
3y - 30 = 90
3y = 90 + 30
3y = 120
y = 40
...
Substitute y into either equation and solve for x
x = 2(40) - 30 = 80 - 30 = 50
...
Solution
x = 50, y = 40
(50,40)
..............................
Have more questions?
HomeworkHelpers@ReadingBoosters.com.
Delighted to help!
-Reading Boosters
Website: www.MyHomeworkAnswers.com